OpenStudy (anonymous):

4x^(3.5) is this a polynomial? and what is the domain range.

OpenStudy (anonymous):

\[4x^(3.5) =4\left( \sqrt{x} \right)^{7}\] \[domain is [0,\infty)\] Because it is defined if x>=0

OpenStudy (anonymous):

\[range is [0,\infty)\]

OpenStudy (anonymous):

okay so the above is the answer to the problem

OpenStudy (anonymous):

it is not a polynomial ?